# A Bayesian joint model of multiple longitudinal and categorical outcomes with application to multiple myeloma using permutation-based variable importance

### v1

@@ Abstract
Joint models have proven to be an effective approach for uncovering potentially hidden connections between various types of outcomes, mainly continuous, time-to-event, and binary. Typically, longitudinal continuous outcomes are characterized by linear mixed-effects models, survival outcomes are described by proportional hazards models, and the link between outcomes are captured by shared random effects. Other modeling variations include generalized linear mixed-effects models for longitudinal data and logistic regression when a binary outcome is present, rather than time until an event of interest. However, in a clinical research setting, one might be interested in modeling the physician’s chosen treatment based on the patient’s medical history in order to identify prognostic factors. In this situation, there are often multiple treatment options, requiring the use of a multiclass classification approach. Inspired by this context, we develop a Bayesian joint model for longitudinal and categorical data. In particular, our motivation comes from a multiple myeloma study, in which biomarkers display nonlinear trajectories that are well captured through bi-exponential submodels, where patient-level information is shared with the categorical submodel. We also present a variable importance strategy for ranking prognostic factors. We apply our proposal and a competing model to the multiple myeloma data, compare the variable importance and inferential results for both models, and illustrate patient-level interpretations using our joint model.
@@ 1 Introduction
In recent years, there has been a growing interest in joint models that simultaneously analyze longitudinal and time-to-event outcomes (Ibrahim et al., 2010; Alsefri et al., 2020). A standard joint model comprises a single longitudinal outcome characterized by a linear mixed-effects submodel, a single survival outcome described by a proportional hazards submodel, and an association structure that connects both submodels (Rizopoulos, 2012). In most cases, this connection involves random effects and complex integrals associated with the survival function, but current literature also presents other alternatives, such as latent classes that stratify heterogeneous subpopulations (Proust-Lima et al., 2014; Andrinopoulou et al., 2020), marginal submodels connected by a copula (Ganjali and Baghfalaki, 2015; Cho et al., 2024), and separate random effects without the need to perform numerical integration (Alvares and Rubio, 2021).
@@
Many authors have extended the standard joint model. For example, some studies modeled multiple longitudinal outcomes and/or competing risks (Hickey et al., 2016; Rué et al., 2017; Rustand et al., 2024), while others incorporated spatial random effects dependence into joint modeling (Martins et al., 2016, 2017; Rappl et al., 2023). There were also proposals that considered categorical longitudinal and survival outcomes (Choi et al., 2015; Chi et al., 2024), two longitudinal submodels, one continuous and one ordinal (Delporte et al., 2024), and joint models of longitudinal continuous and binary data (Wang et al., 2000; Horrocks and van Den Heuvel, 2009; Lu et al., 2016; Zhou et al., 2023). For the latter, to the best of our knowledge, there are currently no available extensions that accommodate more than two categories. This extension introduces at least two additional complexities: (i) more parameters to be estimated, which requires more data, and (ii) imbalanced categories, leading to poor predictive performance (Kolo, 2010).
@@
To fill this gap in the literature, we developed a Bayesian joint model for longitudinal and categorical outcomes. Our proposal is motivated by a retrospective cohort study from a real-world database of patients who were diagnosed with multiple myeloma in the US between January 2015 and February 2022. In this application, there are two nonlinear longitudinal biomarkers that are modeled through bi-exponential submodels (Claret et al., 2009; Thai et al., 2022). The bi-exponential model includes three components (baseline, growth rate, and decay rate) that summarize the temporal dynamics of each biomarker and potentially help to explain the physician’s chosen treatment. Additionally, we considered three treatment options which can be described using a (nominal) categorical submodel (Agresti, 2013). The longitudinal and categorical submodels are connected by sharing the patient-level baseline, growth rate, and decay rate parameters. The main research question here is what are the key factors supporting the physician’s chosen treatment. This is motivated by the clinical knowledge that there is a set of factors contributing to treatment choice, but very rarely are there quantifications of the relative contribution of each of them.
@@
The adequacy of the bi-exponential submodels is verified through an analysis of individual weighted residuals (Desmée et al., 2017), while performance of the categorical submodel is evaluated using metrics for multiclass classification (Grandini et al., 2020). Moreover, as a complement to the posterior inference, we present a model-agnostic variable importance approach with the aim of improving the transparency, trustability, and interpretability of the results (Casalicchio et al., 2019).
@@ 2 Multiple myeloma data
Multiple myeloma (MM) is a type of blood cancer that affects plasma cells. These cancerous plasma cells multiply uncontrollably, eventually crowding out healthy blood cells in the bone marrow, and secrete a monoclonal protein/paraprotein, also known as M-protein (van de Donk et al., 2021). The main MM symptoms are bone pain, weakness, fatigue, frequent infections, and easily broken bones, as well as potentially causing kidney and immune system problems (Kumar et al., 2017). MM treatments may include targeted therapy e.g. proteasome inhibitors and immunomodulatory agents, chemotherapy, stem cell transplant, and immune-based therapies including T-cell engaging antibodies and CAR-T (Rajkumar and Kumar, 2020). While there is currently no cure for MM, the advent of new treatments has led to an increase of progression free survival and overall survival (Gulla and Anderson, 2020).
@@ 2.3 Treatment of continuous baseline covariates
Most continuous covariates contain missing data (10%-56%). In order to address this issue, we apply a log transformation to reduce asymmetry, a standardization (z-score) so that their scales are similar, and a mean imputation to replace missing values. Supplementary Appendix A shows a descriptive summary of these covariates before and after being treated.
@@ 3.1 Longitudinal submodels
To describe the patient-level time evolution of M-spike ($k=1$) and FLC ($k=2$) biomarkers, we consider a bi-exponential model (Stein et al., 2008), which has already been successfully applied in an MM framework (Alvares et al., 2024) and is described as follows:
@@ 3.3 Prior elicitation
We consider default priors for parameters associated with biomarkers ($k=1,2$) and category outcomes ($j=1,2$), but we acknowledge the possibility of using alternative priors. In particular, we assume independent and weakly informative marginal prior distributions. For the bi-exponential submodel (1): population parameters, $\theta_{1k}$, $\theta_{2k}$, and $\theta_{3k}$, follow Normal($0,10^{2}$) prior distributions; residual error variance, $\sigma_{k}^{2}$, follows a half-Cauchy($0,5$) prior distribution; and random effects variance-covariance matrix, ${\bm{\Omega}}_{k}$, follows an inverse-Wishart($\mathbf{I}_{3},4$) prior distribution, where $\mathbf{I}_{3}$ represents a $3\times 3$ identity matrix. For the categorical submodel: regression coefficients, ${\bm{\beta}}_{j}$, and association parameters, $\alpha_{1kj}$, $\alpha_{2kj}$, and $\alpha_{3kj}$, follow Normal($0,10^{2}$) prior distributions.
@@ 4 Variable importance
Variable importance (VI) is a widely known tool in machine learning (Casalicchio et al., 2019), but it has not been widely adopted in statistical modeling. The concept of VI is to measure the importance of covariates by their ability to improve a prediction model’s accuracy. VI and other interpretability techniques can be divided into two groups: model-specific and model-agnostic. Model-specific techniques are based on assumptions imposed by a specific model, whereas model-agnostic ones do not rely on a predefined structure and can be applied to various problems/models without needing extensive customization or fine-tuning (Chen et al., 2023). In this work, we focus on the model-agnostic approach due to its versatility and easy adaptation to different contexts (Ribeiro et al., 2016). In particular, we use the permutation-based VI strategy proposed by Fisher et al. (2019).
@@
The choice of performance metric is an arbitrary decision and may be conditioned on the type of model analyzed. In this work, we used the widely applicable information criterion (WAIC) (Watanabe, 2010), but the VI ranking was identical when applying the leave-one-out cross-validation (Vehtari et al., 2017). Furthermore, due to randomness, calculating VI more than once might greatly vary the results. This problem is mitigated by repeating Algorithm 1 several times and averaging the VI score of each variable. Such a strategy also allows for evaluating the variability of VI scores, for example, by using their minimum and maximum values among the repetitions of the same variable. Finally, it is worth mentioning that after each permutation the model does not require refitting; we only recompute the performance metric.
@@ 5 MM data analysis
Our Bayesian joint model was implemented in Stan using the rstan package version 2.26.23 (Stan Development Team, 2023) from the R language version 4.3.1 (R Core Team, 2023). All codes are publicly available at www.github.com/daniloalvares/BJM-MBiExp-Categ. We set three Markov chains with 4000 posterior samples after 1000 warm-up iterations. Under this configuration, posterior samples achieved convergence (Gelman-Rubin statistic, R-hat $<$ 1.05) and efficiency (effective sample size, neff $>$ 100) (Vehtari et al., 2021).
@@ 6 Discussion
Evaluation of model fit was performed by residual analysis and multiclass classification metrics. IWRES showed good suitability of the bi-exponential submodel to fit the longitudinal trajectories of M-spike and FLC biomarkers, similar to the results obtained by Alvares et al. (2024). The classification metrics also proved to be satisfactory for a three-class problem, mainly when contrasted with the competing model and a random classifier (see Table 1). VI ranking and inference were concise in pointing out the latent characteristics of biomarker trajectories as relevant prognostic factors for choosing treatment. However, it is also important to remember that VI is based on a predictive performance metric, so some differences between VI ranking and inferential significance are expected. Still, VI provides useful and complementary information to the effect sizes, e.g. it proposes a clear ranking of the contribution of each variable to the treatment decision. It is worth mentioning that the treatment decision is a medical choice that may take into account relevant clinical parameters not available in this study, such as comorbidities, fitness, treatment toxicity, patient preference, among others. Finally, we illustrated the use of our joint model for two random patients, showing the longitudinal fit of their biomarkers, and their predicted probabilities for each treatment option.
In conclusion, our study has filled a gap in the modeling literature, shed light on topics related to the clinical practice of MM patients, and connected a popular machine learning strategy to a joint model. As future work, we envision several extensions of the proposed model. For example, in a high-dimensional longitudinal biomarker framework, the robustness to identify significant association parameters is reduced due to potentially strong correlations among biomarkers, but this problem can be overcome using regularization techniques (Andrinopoulou and Rizopoulos, 2016; Sun and Basu, 2024). In the presence of heterogeneous subgroups, we recommend adapting our proposal to joint latent class models (Proust-Lima et al., 2014; Andrinopoulou et al., 2020). In some scenarios, researchers may be interested in studying the causal effect of treatment choice, requiring a causal inference approach (Rizopoulos et al., 2024). Furthermore, the probability for each treatment can be dynamically updated as new longitudinal patient information becomes available (Rizopoulos, 2011; Andrinopoulou et al., 2017). Finally, machine learning algorithms can be proposed for each submodel, such as ensembling techniques for longitudinal data (Hu and Szymczak, 2023) and boosting methods for multiclass classification (Tanha et al., 2020).
@@ Acknowledgements
D.A. and J.K.B. were supported by the U.K. Medical Research Council grant MC_UU_00002/5 and the collaboration grant jointly funded by Roche and the University of Cambridge. We thank the following colleagues for helpful discussions during the research and review stage of this work: Vallari Shah, Mellissa Williamson, Sarwar Mozumder, Madlaina Breuleux, Pascal Chanu, and Chris Harbron. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.
@@ Appendix C. Joint model applied to MM data
Figure 2: Individual weighted residuals (IWRES) from the bi-exponential submodel for each biomarker.

### v2

@@ Abstract
Joint models have proven to be an effective approach for uncovering potentially hidden connections between various types of outcomes, mainly continuous, time-to-event, and binary. Typically, longitudinal continuous outcomes are characterized by linear mixed-effects models, survival outcomes are described by proportional hazards models, and the link between outcomes are captured by shared random effects. Other modeling variations include generalized linear mixed-effects models for longitudinal data and logistic regression when a binary outcome is present, rather than time until an event of interest. However, in a clinical research setting, one might be interested in modeling the physician’s chosen treatment based on the patient’s medical history to identify prognostic factors. In this situation, there are often multiple treatment options, requiring the use of a multiclass classification approach. Inspired by this context, we develop a Bayesian joint model for longitudinal and categorical data. In particular, our motivation comes from a multiple myeloma study, in which biomarkers display nonlinear trajectories that are well captured through bi-exponential submodels, where patient-level information is shared with the categorical submodel. We also present a variable importance strategy to rank prognostic factors. We apply our proposal and a competing model to the multiple myeloma data, compare the variable importance and inferential results for both models, and illustrate patient-level interpretations using our joint model.
@@ 1 Introduction
In recent years, there has been a growing interest in joint models that simultaneously analyze longitudinal and time-to-event outcomes (Ibrahim et al., 2010; Alsefri et al., 2020). A standard joint model comprises a single longitudinal outcome characterized by a linear mixed-effects submodel, a single survival outcome described by a proportional hazards submodel, and an association structure that connects both submodels (Rizopoulos, 2012). In most cases, this connection involves random effects and complex integrals associated with the survival function, but the current literature also presents other alternatives, such as latent classes that stratify heterogeneous subpopulations (Proust-Lima et al., 2014; Andrinopoulou et al., 2020), marginal submodels connected by a copula (Ganjali and Baghfalaki, 2015; Cho et al., 2024), and separate random effects without the need to perform numerical integration (Alvares and Rubio, 2021).
@@
Many authors have extended the standard joint model. For example, some studies modeled multiple longitudinal outcomes and/or competing risks (Hickey et al., 2016; Rué et al., 2017; Rustand et al., 2024), while others incorporated the dependence of spatial random effects into the joint modeling (Martins et al., 2016, 2017; Rappl et al., 2023). There were also proposals that considered categorical longitudinal and survival outcomes (Choi et al., 2015; Chi et al., 2024), two longitudinal submodels, one continuous and one ordinal (Delporte et al., 2024), and joint models of longitudinal continuous and binary data (Wang et al., 2000; Horrocks and van Den Heuvel, 2009; Lu et al., 2016; Zhou et al., 2023). For the latter, to the best of our knowledge, there are currently no extensions available that accommodate more than two categories. This extension introduces at least two additional complexities: (i) more parameters to be estimated, which requires more data, and (ii) imbalanced categories, leading to poor predictive performance (Kolo, 2010).
@@
To fill this gap in the literature, we developed a Bayesian joint model for longitudinal and categorical outcomes. Our proposal is motivated by a retrospective cohort study from a real-world database of patients diagnosed with multiple myeloma in the United States between January 2015 and February 2022. In this application, there are two nonlinear longitudinal biomarkers that are modeled through bi-exponential submodels (Claret et al., 2009; Thai et al., 2022). The bi-exponential model includes three components (baseline, growth rate, and decay rate) that summarize the temporal dynamics of each biomarker and potentially help to explain the physician’s chosen treatment. Furthermore, we considered three treatment options that can be described using a (nominal) categorical submodel (Agresti, 2013). The longitudinal and categorical submodels are connected by sharing the patient-level baseline, growth rate, and decay rate parameters. The main research question here is what are the key factors supporting the physician’s chosen treatment. This is motivated by the clinical knowledge that there is a set of factors contributing to treatment choice, but very rarely are there quantifications of the relative contribution of each of them.
@@
The adequacy of the bi-exponential submodels is verified through an analysis of individual weighted residuals (Desmée et al., 2017), while performance of the categorical submodel is evaluated using metrics for multiclass classification (Grandini et al., 2020). Moreover, as a complement to posterior inference, we present a model-agnostic variable importance approach with the aim of improving the transparency, trustability, and interpretability of the results (Casalicchio et al., 2019).
@@ 2 Multiple myeloma data
Multiple myeloma (MM) is a type of blood cancer that affects plasma cells. These cancerous plasma cells multiply uncontrollably, eventually crowding out healthy blood cells in the bone marrow, and secrete a monoclonal protein/paraprotein, also known as M-protein (van de Donk et al., 2021). The main MM symptoms are bone pain, weakness, fatigue, frequent infections, and easily broken bones, as well as potentially causing kidney and immune system problems (Kumar et al., 2017). MM treatments may include targeted therapy, e.g. proteasome inhibitors and immunomodulatory agents, chemotherapy, stem cell transplant, and immune-based therapies including T-cell engaging antibodies and CAR-T (Rajkumar and Kumar, 2020). While there is currently no cure for MM, the advent of new treatments has led to an increase of progression free survival and overall survival (Gulla and Anderson, 2020).
@@ 2.3 Treatment of continuous baseline covariates
Most continuous covariates contain missing data (10%-56%). In order to address this issue, we apply a log transformation to reduce asymmetry, a standardization (z-score) so that their scales are similar, and a mean imputation to replace missing values. Supplementary Appendix A shows a descriptive summary of these covariates before and after being treated. It is worth mentioning that we also performed the analyses without imputation and, although the sample size was greatly reduced, the conclusions were consistent with those using imputation (see Section 5).
@@ 3.1 Longitudinal submodels
To describe the patient-level time evolution of M-spike ($k=1$) and FLC ($k=2$) biomarkers, we consider a bi-exponential model (Stein et al., 2008), which has already been successfully applied in an MM framework (Alvares et al., 2025) and is described as follows:
@@
It is worth noting that we assumed that the M-spike and FLC trajectories are independent. Mathematically, this assumption was made to overcome convergence issues associated with a common unstructured variance-covariance matrix for all random effects. However, in medical practice, these biomarkers are expected to be complementary but not necessarily correlated (Tacchetti et al., 2017; Gran et al., 2021), which is aligned with what we assumed.
@@ 3.3 Prior elicitation
We consider standard priors for parameters associated with biomarkers ($k=1,2$) and category outcomes ($j=1,2$), but we acknowledge the possibility of using alternative priors. In particular, we assume independent and weakly informative marginal prior distributions. For the bi-exponential submodel (1): population parameters, $\theta_{1k}$, $\theta_{2k}$, and $\theta_{3k}$, follow Normal($0,10^{2}$) prior distributions; residual error variance, $\sigma_{k}^{2}$, follows a half-Cauchy($0,5$) prior distribution; and random effects variance-covariance matrix, ${\bm{\Omega}}_{k}$, follows an inverse-Wishart($\mathbf{I}_{3},4$) prior distribution, which is a minimally informative specification (Schuurman et al., 2016), where $\mathbf{I}_{3}$ represents a $3\times 3$ identity matrix. For the categorical submodel: regression coefficients, ${\bm{\beta}}_{j}$, and association parameters, $\alpha_{1kj}$, $\alpha_{2kj}$, and $\alpha_{3kj}$, follow Normal($0,10^{2}$) prior distributions. As a sensitivity analysis, we also set Normal($0,100^{2}$) and half-Cauchy($0,25$) prior distributions in order to make them less informative. The results were equivalent in terms of posterior distributions obtained, differing only in computational time, so we concluded that our initial choice is weakly informative and speeds up the processing time to achieve convergence of Markov chains.
@@ 4 Variable importance
Variable importance (VI) is a widely known tool in machine learning (Casalicchio et al., 2019), but it has not been widely adopted in statistical modeling. The concept of VI is to measure the importance of covariates by their ability to improve the prediction model’s accuracy. VI and other interpretability techniques can be divided into two groups: model-specific and model-agnostic. Model-specific techniques are based on assumptions imposed by a specific model, whereas model-agnostic ones do not rely on a predefined structure and can be applied to various problems/models without need for extensive customization or fine-tuning (Chen et al., 2023). In this work, we focus on the model-agnostic approach due to its versatility and easy adaptation to different contexts (Ribeiro et al., 2016). In particular, we use the permutation-based VI strategy proposed by Fisher et al. (2019).
@@
The choice of performance metric is an arbitrary decision and may be conditioned on the type of model analyzed. In this work, we used the widely applicable information criterion (WAIC) (Watanabe, 2010), but the VI ranking was identical when applying the leave-one-out cross-validation (Vehtari et al., 2017). Furthermore, due to randomness, calculating VI more than once might greatly vary the results. This problem is mitigated by repeating Algorithm 1 several times and averaging the VI score of each variable. This strategy also allows us to evaluate the variability of VI scores, for example, by using their minimum and maximum values among the repetitions of the same variable. Finally, it is worth mentioning that after each permutation the model does not require refitting; we only recompute the performance metric.
@@ 5 MM data analysis
Our Bayesian joint model was implemented in Stan using the rstan package version 2.26.23 (Stan Development Team, 2023) from the R language version 4.3.1 (R Core Team, 2023). All codes are publicly available at www.github.com/daniloalvares/BJM-MBiExp-Categ. We set three Markov chains with 4000 posterior samples after 1000 warm-up iterations. Under this configuration, posterior samples achieved convergence (Gelman-Rubin statistic, R-hat $<$ 1.05) and efficiency (effective sample size, neff $>$ 500) (Vehtari et al., 2021).
@@ 6 Discussion
Evaluation of model fit was performed by residual analysis and multiclass classification metrics. IWRES showed good suitability of the bi-exponential submodel to fit the longitudinal trajectories of M-spike and FLC biomarkers, similar to the results obtained by Alvares et al. (2025). The classification metrics also proved to be satisfactory for a three-class problem, mainly when contrasted with the competing model and a random classifier (see Table 1). VI ranking and inference were concise in pointing out the latent characteristics of biomarker trajectories as relevant prognostic factors for choosing treatment. However, it is also important to remember that VI is based on a predictive performance metric, so some differences between the VI ranking and the inferential significance are expected. Still, VI provides useful and complementary information to the effect sizes, e.g., it proposes a clear ranking of the contribution of each variable to the treatment decision. It should be mentioned that the treatment decision is a medical choice that may take into account relevant clinical parameters not available in this study, such as comorbidities, fitness, treatment toxicity, patient preference, among others. Finally, we illustrated the use of our joint model for two random patients, showing the longitudinal fit of their biomarkers, and their predicted probabilities for each treatment option.
In conclusion, our study has filled a gap in the modeling literature, shed light on topics related to the clinical practice of MM patients, and connected a popular machine learning strategy to a joint model. In future work, we envision several extensions of the proposed model. For example, in a high-dimensional longitudinal biomarker framework, the robustness to identify significant association parameters is reduced due to potentially strong correlations among biomarkers, but this problem can be overcome using regularization techniques (Andrinopoulou and Rizopoulos, 2016; Sun and Basu, 2024). We chose a VI strategy as a method for ranking the variables, but the Bayesian framework also provides us with spike-and-slab priors as an alternative approach (Islam et al., 2024). In the presence of heterogeneous subgroups, we recommend adapting our proposal to joint latent class models (Proust-Lima et al., 2014; Andrinopoulou et al., 2020). In some scenarios, researchers may be interested in studying the causal effect of treatment choice, which requires a causal inference approach (Rizopoulos et al., 2024). Furthermore, the probability for each treatment can be dynamically updated as new longitudinal patient information becomes available (Rizopoulos, 2011; Andrinopoulou et al., 2017). Finally, machine learning algorithms can be proposed for each submodel, such as ensembling techniques for longitudinal data (Hu and Szymczak, 2023) and boosting methods for multiclass classification (Tanha et al., 2020).
@@ Acknowledgements
D.A. and J.K.B. were supported by the U.K. Medical Research Council grant MC_UU_00002/5 and the collaboration grant jointly funded by Roche and the University of Cambridge. The authors thank the following colleagues for helpful discussions during the research and review stage of this work: Vallari Shah, Mellissa Williamson, Sarwar Mozumder, Madlaina Breuleux, Pascal Chanu, and Chris Harbron. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.
@@ Appendix C. Joint model applied to MM data
Figure 2: Individual weighted residuals (IWRES) from the bi-exponential submodel for each biomarker. For suitable fit, IWRES should approximately follow a Normal distribution centered at zero (Kerioui et al., 2022).